OpenStudy (anonymous):

The area of a triangle is 16 square units. The base of the triangle is x+4. The height of the triangle is x. Find x.

OpenStudy (anonymous):

what is the formula for finding the area of a triangle?

OpenStudy (anonymous):

A = 1/2bh

OpenStudy (anonymous):

so you plug in x+4 for b and x for h and set it all equal to 16 and solve for x

OpenStudy (anonymous):

still don't know the answer.

OpenStudy (anonymous):

can you set up the equation?

OpenStudy (anonymous):

Im so lost ... Sorry

OpenStudy (anonymous):

you know the area is 16, that is the A so A = 1/2 b h becomes 16 = 1/2 b h now you know b is x+4 and h is x so plug them in the same way

OpenStudy (anonymous):

so x equals 8

OpenStudy (anonymous):

why does x equal 8? I don't see it

OpenStudy (anonymous):

You said A = 1/2 bh so if I SAY 1/2 * 8*4 that equals 16

OpenStudy (anonymous):

if x = 8 then what is x+4 ? the equation you need to solve is: 16 = 1/2 (x) (x+4)

OpenStudy (anonymous):

um okay Thanks anyways !

OpenStudy (anonymous):

all right, good luck with it sorry if I confused you, but x is not 8 here